How many moles of O2 are needed to react with 2.83 moles 


1
Expert's answer
2022-03-19T02:45:00-0400

n(H2O) = 2.83 moles.


Solution:

The equation of a chemical reaction:

4NO2(g) + 6H2O(g) → 7O2(g) + 4NH3(g)

According to the reaction equation: n(NO2)/4 = n(H2O)/6.

n(NO2) = 4*n(H2O)/6;

n(NO2) = (4 * 2.83 moles) / 6 = 1.8867 moles = 1.89 moles

n(NO2) = 1.89 moles.


Answer: 1.89 moles of nitrogen dioxide (NO2).